OpenStudy (anonymous):

find dx/dy x^2 + xy - y^3 = xy^2

OpenStudy (anonymous):

\[x^2 + xy - y^3 = xy^2\] And please help me in showing steps trying to get concept

OpenStudy (apoorvk):

dx/dy sure?? and not dy/dx??

OpenStudy (anonymous):

you are right.... I typed it wrong sorry :)

OpenStudy (anonymous):

2x+x(dy/dx) +y -3y^2(dy/dx)=2xy(dy/dx)+y^2 Thats it :)

OpenStudy (apoorvk):

okay... hmm differentiate both sides. apply chain rule for diferentiate 'y', since it is a function of x. and use product for diffrentiating xy and xy^2 type if terms. try once.

OpenStudy (anonymous):

I got \[(2x - y^2 + y) / (2xy - x +3y^2)\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes Thats Right :)

OpenStudy (anonymous):

great! thanks just checking myself

OpenStudy (anonymous):

No issues..:)